OpenStudy (sphott51):

simplify (x^2 -9) / (x^2 +x -6) The answer is (x-3) / (x-2 Don't know how to get there, needs to show the work

OpenStudy (loser66):

\(a^2-b^2 =(a+b)(a-b)\) for numerator, your a is x, your b is 3 (since 3^2=9),

OpenStudy (loser66):

for denominator, just factor it to get (x+3)(x-2)

OpenStudy (loser66):

combine them and cancel, you get the answer.

OpenStudy (sphott51):

Thanks